OpenStudy (anonymous):

SOS!! x + y = 6 x - y = 8 Solve the system of equations.

OpenStudy (solomonzelman):

Can you see that y is a negative number ?

OpenStudy (acxbox22):

add the 2 equations, eliminating the y solve for x plug x value into any equation to get y value

OpenStudy (solomonzelman):

7+(-1)=6 7-(-1)=8

OpenStudy (solomonzelman):

pure logic ;)

OpenStudy (texaschic101):

just add them and you will see that the y's cancel out...and just solve for x. Once x is found, sub that back into either equation and solve for y

OpenStudy (anonymous):

Don't be sassy Soloman

OpenStudy (texaschic101):

lol

OpenStudy (anonymous):

I tried, I need the answer because mine are wrong!

OpenStudy (acxbox22):

@SolomonZelman gave you the answer

OpenStudy (acxbox22):

look at the 3rd post in this thread

OpenStudy (texaschic101):

x + y = 6 x - y = 8 -----------add 2x = 14 x = 7 now sub 7 in for x in either equation to find y x + y = 6 7 + y = 6 y = 6 - 7 y = -1 check.. x - y = 8 7 - (-1) = 8 7 + 1 = 8 8 = 8 (correct) x = 7 and y = -1 any questions ?

OpenStudy (anonymous):

Thank you :)

OpenStudy (triciaal):

follow the directions given by @acxbox22

OpenStudy (texaschic101):

:)